OpenStudy (anonymous):

Solve for x: (4^x)+6(4^-x)=5 Medal for Answer

OpenStudy (anonymous):

is it \[ 4^x+6(4^{-x})=5\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hey u there?

OpenStudy (anonymous):

So lets apply ln and see if we make sense \[ \ln (4^x)+ \ln (6*4^{-x})=\ln (5)\] \[x \ln (4) +ln(6)-x \ln (4) = \ln (5)\] Ya I am here. Just trying to see how to continue

OpenStudy (anonymous):

Actually I gotta go home so I will be back in like 1/2 hr

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

makes sense now